OpenStudy (anonymous):

Seven more than the average of 4 numbers is 52. Find the sum of the four numbers.

OpenStudy (anonymous):

(a+b+c+d)/4 + 7 = 52 (a+b+c+d)/4=52-7=45 (a+b+c+d)=45*4=?